If one month there is an increase of 50% and the next there is a decrease of 25% how much was the increase or decrease.
Answer by stanbon(48535)

If one month there is an increase of 50% and the next there is a decrease of 25% how much was the increase or decrease.
---------------
Start with "x".
After one month you have x + 0.50x = 1.5x
After the next you have 1.5x-0.25(1.5x) = 1.5x - 0.3750x = 1.125x
The increase was 0.125x = 12.5%
=================
Cheers,
Stan H.